How long will it take $5,000 investment to grow to $6,500 at an annual rate of 7% compounded semi-annually
Answer #1

PV =5000
FV =6500
Rate Semi annually =7%/2 =3.5%
Using Financial Calculator to calculate Number of Periods
I/Y =3.5%;PV=5000;FV=-6500;CPT N =7.6266

Number of years =7.6266/2 =3.81 years